OpenStudy (anonymous):

What is the d value of the following sequence? 0,1/2,1,3/2,2

OpenStudy (anonymous):

c.d=t2-t1

OpenStudy (anonymous):

??

OpenStudy (anonymous):

im sorry..im confuessed

OpenStudy (anonymous):

\[c.d.=\frac{ 1 }{ 2 }-0=\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

what the c for?